Evening Shift Summary 07/13

Renata Zukanovich-Funchal (zukanov@fnal.gov)
Wed, 14 Jul 1999 00:13:09 -0500

Shift Summary for evening shift, 07/13/99

On Shift: E. Blucher, R. Zukanovich Funchal

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 1 h

Accounting of lost time: Tevatron scheduled down time and ground fault
repair (MI Quad replacement).

========================================================================
Runs ended this shift:

Run 13893 LASER_SCAN
Run 13896 LASER_SCAN

Tapes brought to FCC this shift: none
========================================================================
Accesses During Shift:

========================================================================
Overview of Shift:
special runs; work on HA shaper/splitter power supply.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 Took over. No beam.

17:30 Problems with trigger, camac. Reloading busy logic
cleared trigger problem. Branch 2 camac problem
cured itself.

13893 18:40 Laser scan starts.

19:17 End of laser run. We take a cosmic run.

19:35 Ktev3 disk full. Onsplit requests QKPE06 to be
mounted. We do so.

13896 22:00 Special LASER_SCAN_HCC changed 1136 HV.

23:30 Search and Secure was done.

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift
- still problems with HA shaper/splitter power supply.